find the radius R of the largest sphere that can be fitted into a cylinder of radius 'r' and height 'h' in terms of 'r' and 'h'.

Grade:10

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

R=radius of sphere

x= the hight of the cylinder

r=radius of cylinder

r^2+x^2=(2R)^2 r^2=-x^2+4R^2

v(x)=pi*r^2*h=pi(-x^2+4R^2)*x=pi*(-x^3…


v'=pi(-3x^2+4R^2)=0 3x^2=4R^2

x^2=4R^2/3

r^2=-x^2+4R^2=-4R^2/3+4R^2=4R^2*2/3=

8R^2/3 r=sqrt(8R^2/3 )=Rsqrt(8/3)

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi